# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
For $n=50$, the $50\times 50$ covariance matrix $\Sigma$ can easily become ill-conditioned. Even if the BPTT-derived estimate for $V_{x\mathbf{Y}}$ is quite accurate, any small residual noise can be significantly amplified when multiplied by $\Sigma^{-1}$ if its condition number is large. This amplification effect is the primary suspect for the observed hedging error.
@@
In conclusion, the $n=50,k=10$ visualizations powerfully demonstrate that Projected PG-DPO (P-PGDPO) can effectively ”break the dimensional barrier” where Baseline PG-DPO fails and Deep BSDE offers only a partial (myopic-focused) solution. The remaining inaccuracies in the hedging component point towards the need for addressing the ill-conditioning of the covariance matrix in very high asset dimensions.
@@ 5 Conclusion
To surmount these obstacles, we introduced the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, with its highly effective Projected PG-DPO (P-PGDPO) variant (Section 3.2). P-PGDPO distinctively leverages Pontryagin’s Maximum Principle (PMP) by first employing backpropagation-through-time (BPTT) to achieve rapidly stabilizing estimates of the Pontryagin costates and their derivatives (e.g., $\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$). These estimates are then analytically projected onto the manifold of PMP-optimal controls via the first-order conditions. This decoupling strategy proved remarkably successful, enabling P-PGDPO to tackle problems of previously prohibitive scales. For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.58\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4). The theoretical underpinnings of P-PGDPO are strengthened by our Policy Gap Bound (Theorem 1), which guarantees near-optimal policies given accurate costate estimation and a small FOC gap.
@@
The practical implications of P-PGDPO are significant. Its demonstrated ability to recover complex, non-linear hedging strategies in high dimensions suggests its potential applicability to large institutional portfolios, such as those involving numerous ETFs or derivative hedging. Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure (Section 3.3).
@@
  • Extending the P-PGDPO framework to incorporate portfolio constraints (e.g., no short-selling, leverage limits) and market frictions like transaction costs, by adapting the PMP conditions accordingly.
@@
  • Validating the framework through rigorous backtesting on historical and simulated market data for non-affine models.
@@ Appendix B Analytic Solutions for Affine Models
$\quad-\left[\mathbf{C}(t)\kappa_{Y}-\frac{1-\gamma}{\gamma}(\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right]\bm{\theta}_{Y}.$ (25)
@@
$\frac{dg(t)}{dt}$ $=(1-\gamma)r$
$\quad-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)$
$\quad-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
@@
$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}$
$\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$ (26)
@@
Appendix C Detailed Proofs for the Policy Gap Bound (Theorem 1) and Supporting Results
This appendix provides the detailed mathematical arguments supporting the Policy Gap Bound stated in Theorem 1 (see Section 3.3). These details include the precise statements of the necessary assumptions (Assumptions 6 and 7), and the full proof sketches for the lemmas, propositions, and Theorem 4 that collectively establish the theoretical justification for the Projected PG-DPO approach. The core result demonstrates that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$.
Objective and Roadmap.
This section provides the mathematical justification for the Projected PG-DPO approach. We establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$. Our proof follows the logical chain:
@@ Appendix C Detailed Proofs for the Policy Gap Bound (Theorem 1) and Supporting Results
The core involves establishing bounds via lemmas and propositions, resolving coupled estimates, and concluding with the main theorem.
1. Preliminaries: PDE, Norms, and Regularity.
We work within the framework established in Section 2. For any admissible policy $(\pi,c)$, the associated sub-value function $V^{\pi,c}$ solves the linear parabolic PDE:
$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\quad V^{\pi,c}(T,z)=g(z),$ (27)
where the operator $\mathcal{L}_{\pi,c}$ and source term $f_{\pi,c}$ depend on the policy $(\pi,c)$. This dependence occurs through specific functions determining the PDE coefficients and the source term. For the purpose of tracking policy influence in our analysis, we conceptually group the key policy-dependent components under the label (27) below :
  • Wealth Drift Influence: Terms multiplying $V_{x}$ dependent on $\pi_{t},C_{t}$.
  • Wealth Diffusion Influence: Term multiplying $V_{xx}$ dependent on $\bm{\pi}_{t}$.
  • Wealth-State Cross-Diffusion Influence: Terms multiplying $V_{xy_{j}}$ dependent on $\bm{\pi}_{t}$.
  • Consumption Source Term: $f_{\pi,c}=e^{-\delta t}U(C_{t})$.
We use standard mixed $L^{q,p}(Q_{T})$ space-time norms over the domain $Q_{T}=(0,T)\times\Omega$, where $\Omega\subseteq\mathbb{R}^{d}$ is the spatial domain, with $p>d+2$ and $q\geq 2$. We define the error norms: $E_{f}=\|f_{\pi^{*},c^{*}}-f_{\pi,c}\|_{L^{q,p}}$, $E_{h}=\|h\|_{L^{q,p}}$ (where $h$ represents the difference in policy influence on PDE coefficients as per (27)), and $E_{v}=\|\nabla V^{\pi^{*},c^{*}}-\nabla V^{\pi,c}\|_{L^{q,p}}$.
2. Standing Assumptions.
Our analysis relies on standard regularity conditions.
Assumption 6 (Parabolic Regularity).
We assume: (a) The maps from policy $(\pi,c)$ to PDE coefficients (drift $b$, diffusion $a$) are Lipschitz continuous with constant $L_{\mathrm{coef}}$. (b) The Hamiltonian $H(t,z,V,X,\pi,c)$ is $C^{2}$ in its arguments; its gradient $\nabla_{(\pi,c)}H$ is Lipschitz continuous in $X:=\nabla V$ with constant $L_{HX}$. (c) The map from costates to controls via PMP FOCs, $F:(\lambda,\nabla\lambda)\mapsto(\pi,c)$, is locally Lipschitz. Standard uniform parabolicity holds: $\xi^{\top}a(t,z,\pi,c)\xi\geq\underline{\nu}\|\xi\|^{2}$ for some $\underline{\nu}>0$. Coefficients are sufficiently smooth for PDE regularity results (e.g., VMO in space, bounded measurable in time).
Assumption 7 (FOC Gap).
The policy $(\pi,c)$ from Stage 1 and its costate $\lambda=\nabla V^{\pi,c}$ satisfy:
$\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon,$
where $\mathcal{R}_{\mathrm{FOC}}:=\nabla_{(\pi,c)}H(t,z,V^{\pi,c},\nabla V^{\pi,c},\pi,c)$.
3. Key Technical Results.
Lemma 2 (FOC–reduction: coefficient–gap bound, full version).
Let Assumptions 67 hold and assume further that
$\exists\,\underline{\Lambda}>0:\qquad-\nabla_{(\pi,c)}^{2}H\;\succeq\;\underline{\Lambda}I_{(n+1)\times(n+1)}\quad\text{for all }(t,z,\pi,c)\text{ in the admissible set }{\mathcal{A}},$ (28)
i.e. the Hamiltonian is uniformly strongly concave in the control variables. Let $h(t,z)$ collect the differences of every coefficient (in drift, diffusion and cross terms) of the linear operator $\mathcal{L}_{\pi,c}$ when $(\pi,c)$ is substituted for the optimal control $(\pi^{*},c^{*})$. Then, for all $(t,z)\in(0,T)\times\Omega$,
$|h(t,z)|\;\leq\;L_{H}\,\bigl{|}\nabla V^{\pi,c}(t,z)-\nabla V^{*}(t,z)\bigr{|}\;+\;C_{\mathrm{FOC}}\,\bigl{|}\mathcal{R}_{\mathrm{FOC}}(t,z)\bigr{|},$ (29)
with finite constants
$L_{H}:=L_{\mathrm{coef}}\,\frac{L_{HX}}{\underline{\Lambda}},\qquad C_{\mathrm{FOC}}:=L_{\mathrm{coef}}\,\frac{1}{\underline{\Lambda}},$ (30)
where $L_{\mathrm{coef}}$ and $L_{HX}$ are as defined in Assumption 6. (Dependence: $L_{H},C_{\mathrm{FOC}}$ depend only on $\underline{\Lambda},L_{\mathrm{coef}},L_{HX}$ and therefore ultimately on the model parameters fixed in Assumption 6.)
@@
First, we define the implicit system $F(\pi,c;V,X):=\nabla_{(\pi,c)}H(t,z,V,X,\pi,c)$, where $X:=\nabla V$. For every fixed state point $(t,z)$, the first–order condition is $F(\pi,c;V,X)=0$. The optimal pair $(\pi^{*},c^{*})$ satisfies $F(\pi^{*},c^{*};V^{*},\nabla V^{*})=0$.
Next, by the uniform strong concavity (28), $\nabla_{(\pi,c)}F=-\nabla_{(\pi,c)}^{2}H$ is invertible on $\mathcal{A}$ and $\bigl{\|}(\nabla_{(\pi,c)}F)^{-1}\bigr{\|}\leq\underline{\Lambda}^{-1}$. Hence, the Implicit–Function Theorem yields a $C^{1}$-map $\Phi:(V,X)\mapsto(\pi,c)$ on a neighbourhood of $(V^{*},X^{*})$ with $\bigl{\|}\nabla\Phi(V,X)\bigr{\|}=\bigl{\|}(\nabla_{(\pi,c)}F)^{-1}\nabla_{(V,X)}F\bigr{\|}\leq\underline{\Lambda}^{-1}\,L_{HX}=:C_{H}$ for all $(V,X)$ in the neighbourhood.
Then, we establish a mean–value bound for the control gap. For the non–optimal pair $(\pi,c)$, we have $(\pi,c)-(\pi^{*},c^{*})=\Phi\!\bigl{(}V^{\pi,c},\nabla V^{\pi,c}\bigr{)}-\Phi\!\bigl{(}V^{*},\nabla V^{*}\bigr{)}$. By the mean–value theorem in Banach spaces, $|(\pi,c)-(\pi^{*},c^{*})|\leq C_{H}\,\bigl{|}\nabla V^{\pi,c}-\nabla V^{*}\bigr{|}+\underline{\Lambda}^{-1}\,\bigl{|}F(\pi,c;V^{\pi,c},\nabla V^{\pi,c})\bigr{|}$. The second term equals $|\mathcal{R}_{\mathrm{FOC}}|$ by Assumption 7.
Finally, we translate this to a coefficient gap. Because the mapping $(\pi,c)\mapsto(b,a)$ (PDE coefficients) is $L_{\mathrm{coef}}$-Lipschitz (Assumption 6(a)), $|h(t,z)|\;\leq\;L_{\mathrm{coef}}\,|(\pi,c)-(\pi^{*},c^{*})|$. Substituting the estimate from the previous paragraph gives $|h|\leq L_{\mathrm{coef}}C_{H}\,|\nabla V-\nabla V^{*}|\;+\;L_{\mathrm{coef}}\underline{\Lambda}^{-1}\,|\mathcal{R}_{\mathrm{FOC}}|$. The stated constants $L_{H},C_{\mathrm{FOC}}$ now follow by definition. ∎
Lemma 3 (Source–gap bound, full version).
Fix $p>d+2,\;q\geq 2$ and let the standing assumptions 67 hold. Assume furthermore that there exist deterministic constants
$0<\underline{c}<\overline{c}<\infty,\qquad 0<\underline{\lambda}<\overline{\lambda}<\infty,$ (31)
such that for every admissible trajectory and for a.e. $t\in[0,T]$
$\underline{c}\;\leq\;c_{t},\;c^{*}_{t}\;\leq\;\overline{c},\quad\underline{\lambda}\;\leq\;\lambda_{t},\;\lambda^{*}_{t}\;\leq\;\overline{\lambda},$
where $\lambda_{t}:=V^{\pi,c}_{x}(t,\cdot)$ and $\lambda^{*}_{t}:=V^{*}_{x}(t,\cdot)$. Let
$E_{f}\;:=\;\bigl{\|}\,e^{-\delta t}\bigl{(}U(c^{*})-U(c)\bigr{)}\bigr{\|}_{L^{q,p}},\qquad E_{v}:=\|\nabla V^{*}-\nabla V^{\pi,c}\|_{L^{q,p}},$
and write $\varepsilon:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}$ for the first–order-condition residual norm. Then
$E_{f}\;\leq\;C_{f}\,\varepsilon\;+\;C_{f}^{\prime}\,E_{v},$ (32)
with explicit finite constants
$C_{f}\;=\;\overline{c}^{-\gamma}\,L_{(U^{\prime})^{-1}},\qquad C_{f}^{\prime}\;=\;L_{(U^{\prime})^{-1}},\qquad L_{(U^{\prime})^{-1}}=\frac{\overline{\lambda}^{\frac{\gamma+1}{\gamma}}}{\gamma}\,\underline{c}^{\gamma}.$ (33)
(All dependence on the model is through $\delta,\gamma,\underline{c},\overline{c},\underline{\lambda},\overline{\lambda}$.)

### v4

@@ 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
For $n=50$, the $50\times 50$ matrix $\Sigma$ is prone to ill-conditioning. Thus, even small residual noise in the BPTT-estimated $V_{x\mathbf{Y}}$ may be significantly amplified by $\Sigma^{-1}$, especially when its condition number is large. This amplification effect is the most likely cause of the observed hedging error.
@@
In conclusion, the $n=50,k=10$ visualizations clearly demonstrate that Projected PG-DPO is capable of “breaking the dimensional barrier,” where Baseline PG-DPO fails and Deep BSDE provides only a partial (myopic-focused) solution. The remaining inaccuracies in the hedging component point to a pressing need for addressing the ill-conditioning of $\Sigma$ in very high asset dimensions—potentially through regularization, shrinkage techniques, or architecture-aware preconditioning.
4.3 Non-affine Asset Return Dynamics
In this section, we extend our numerical investigation to a setting where the asset prices evolve according to a non-affine process. We assume that the state factors state factors, $\mathbf{Y}_{t}=[Y_{1,t},\dots,Y_{k,t}]^{T}$, follow the same multivariate Ornstein-Uhlenbeck (OU) process as described in Section 2.1. The expected returns of risky assets, $\bm{\mu}_{S}(t,\mathbf{Y}_{t})$, are given by
$\bm{\mu}_{S}(t,\mathbf{Y}_{t})=\text{diag}(\bm{\sigma})(\bm{\alpha}\mathbf{Y}_{\beta,t})+r\mathbf{1}=\text{diag}(\bm{\sigma})\left(\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})\right)+r\mathbf{1},$
where $\bm{\alpha}$ is a $d\times k$ matrix of factor loadings and the multiplication and addition in $\bm{\alpha}(\mathbf{Y}_{t}+\bm{\beta}\mathbf{Y}_{t}^{2})$ is understood as element-wise operations. The vector $\bm{\beta}$ controls the strength and direction of the non-linear impact of $\mathbf{Y}_{t}$ on asset returns. Its overall magnitude is parameterized by $\beta_{\text{norm}}=||\bm{\beta}||_{2}$, and for our experiments, $\bm{\beta}$ is generated as $\bm{\beta}=\beta_{\text{norm}}\cdot\mathbf{u}$, where $\mathbf{u}$ is a randomly generated unit vector (with a fixed seed for consistency). When $\beta_{\text{norm}}=0$, $\bm{\beta}=\mathbf{0}$, which implies $\mathbf{Y}_{\beta,t}=\mathbf{Y}_{t}$. In this case, the model for $\bm{\mu}_{S}$ becomes affine in $\mathbf{Y}_{t}$, and the overall portfolio optimization problem reduces to the affine setting analyzed in Sections 4.1 and 4.2.
The primary objective of this analysis is twofold: first, to validate the solutions obtained by our Projected PG-DPO (P-PGDPO) framework for this model by examining their behavior as $\beta_{\text{norm}}\to 0$, where we expect convergence to the known affine solution. Second, we investigate how the optimal policy systematically changes as the strength of this non-linear component in the risk premium, dictated by $\beta_{\text{norm}}$, increases. All experiments in this subsection are conducted for the most challenging setting, $n=50$ assets and $k=10$ state factors, utilizing the P-PGDPO method.
To quantify the deviation of the P-PGDPO policy in this non-affine setting from the affine benchmark solution as $\beta_{\text{norm}}$ changes, Table 5 presents the Root Mean Squared Error (RMSE) of the P-PGDPO learned policy, $\bm{\pi}(\beta_{\text{norm}})$, relative to the affine analytical solution ($\bm{\pi}(\beta_{\text{norm}}=0)$). For the results presented in this table, all P-PGDPO models were trained for 10,000 iterations. In the non-affine case ($\beta_{\text{norm}}>0$), where an analytical solution to the control problem is unavailable, this fixed number of iterations is used for evaluation, as opposed to selecting the number of iterations based on the minimum error against a known benchmark in earlier sections.
Table 5: Policy RMSE of P-PGDPO for the model with non-affine risk premia vs. affine benchmark ($n=50,k=10$). Results reported after 10,000 training iterations.
$\beta_{\text{norm}}$ Min. Policy RMSE (P-PGDPO)
0.0 0.04701
0.1 0.05686
0.5 0.06755
1.0 0.11469
2.0 0.20120
3.0 0.27530
4.0 0.33570
The results in Table 5 clearly demonstrate two key points. First, for $\beta_{\text{norm}}=0.0$, the P-PGDPO method achieves a low RMSE of $0.04701$. This is consistent with its strong performance in the affine setting (reported in Section 4.1) and indicates an accurate approximation of the affine analytical solution, as expected when the non-linear component vanishes. Second, as $\beta_{\text{norm}}$ increases from $0.1$ to $4.0$, the difference of the P-PGDPO policy from the fixed affine benchmark, measured by RMSE, monotonically increases. This trend signifies that the optimal policy systematically diverges from the affine optimal policy as the magnitude of the non-linear component grows. The consistent increase suggests that P-PGDPO is stably adapting the policy to this increasing non-linearity in the market price of risk, rather than producing erratic or unstable solutions.
(a) Analytical solution ($\beta_{\text{norm}}=0$).
(b) P-PGDPO policy ($\beta_{\text{norm}}=0.0$).
(c) P-PGDPO policy ($\beta_{\text{norm}}=0.1$).
(d) P-PGDPO policy ($\beta_{\text{norm}}=1.0$).
(e) P-PGDPO policy ($\beta_{\text{norm}}=2.0$).
Figure 5: Visual comparison of optimal policies for the first asset ($\pi_{1}$) against the first state factor ($Y_{1}$) and time-to-maturity ($T-t$) in the $n=50,k=10$ setting. (a) shows the analytical solution for the affine model ($\beta_{\text{norm}}=0$). (b)-(e) display P-PGDPO learned policies for models with increasing strength of non-linearity in asset risk premia, corresponding to $\beta_{\text{norm}}$ values of $0.0,0.1,1.0,2.0$, respectively. Note the systematic change in policy structure and value ranges (indicated by colorbars) as $\beta_{\text{norm}}$ increases.
To complement these quantitative findings, Figure 5 provides a visual comparison of the learned policies. Figure 5(a) displays the analytical solution for the affine model ($\beta_{\text{norm}}=0$). Figures 5(b)–5(e) show the P-PGDPO learned policies for models where $\beta_{\text{norm}}=0.0,0.1,1.0,$ and $2.0$, respectively. These plots depict the policy for the first asset ($\pi_{1}$) as a function of the first state factor ($Y_{1}$) and time-to-maturity ($T-t$), holding other state factors at their long-term means.
As observed in Figure 5(b), the P-PGDPO policy for $\beta_{\text{norm}}=0.0$ closely matches the analytical affine solution shown in Figure 5(a), visually confirming the low RMSE reported in Table 5. As $\beta_{\text{norm}}$ increases to $0.1$ (Figure 5(c)), $1.0$ (Figure 5(d)), and $2.0$ (Figure 5(e)), a clear and systematic transformation of the policy surface is evident. The policy values tend to increase in certain regions (indicated by the shift towards warmer colors and the expanding range of the colorbars), and the overall shape of the policy function changes. This visual progression corroborates the quantitative results presented in Table 5, illustrating that the P-PGDPO framework captures the impact of the non-linear term in the risk premium in a consistent manner. Since the non-affine component $\bm{\beta}\,\mathbf{Y}_{t}^{2}$ directly alters how state variables $\mathbf{Y}_{t}$ influence the investment opportunity set (via $\bm{\mu}_{S}$), these changes in the overall policy are likely significantly driven by adjustments to the intertemporal hedging component.
In summary, the combined quantitative and visual results from this analysis strongly suggest that the P-PGDPO method provides robust and consistent solutions for portfolio optimization problems where the asset prices follow general non-affine dynamics. The learned policies correctly converge to the affine solution as the specific non-linear component vanishes ($\beta_{\text{norm}}\to 0$), and they exhibit a systematic and stable response to increasing levels of this non-linearity. This provides confidence in applying PG-DPO to more complex financial models with non-affine dynamics where analytical solutions are typically unavailable.
@@ 5 Conclusion
To surmount these obstacles, we introduced the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, with its highly effective Projected PG-DPO (P-PGDPO) variant (Section 3.2). P-PGDPO distinctively leverages Pontryagin’s Maximum Principle (PMP) by first employing backpropagation-through-time (BPTT) to achieve rapidly stabilizing estimates of the Pontryagin costates and their derivatives (e.g., $\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$). These estimates are then analytically projected onto the manifold of PMP-optimal controls via the first-order conditions. This decoupling strategy proved remarkably successful, enabling P-PGDPO to tackle problems of previously prohibitive scales. For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.52\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4). The theoretical underpinnings of P-PGDPO are strengthened by our Policy Gap Bound (Theorem 1), which guarantees near-optimal policies given accurate costate estimation and a small FOC gap.
@@
The practical implications of P-PGDPO are significant. Its demonstrated ability to recover complex, non-linear hedging strategies in high dimensions suggests its potential applicability to large institutional portfolios, such as those involving numerous ETFs or derivative hedging. Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure. Indeed, the robustness of the P-PGDPO framework is further underscored in a companion paper, where we successfully apply it to large-scale portfolio optimization under portfolio constraints (Huh et al., , 2025).
@@
  • Extending the P-PGDPO framework to incorporate market frictions such as transaction costs and taxes, by adapting the PMP conditions accordingly.
@@
  • Validating the framework through rigorous backtesting on historical and simulated market data.
@@ Appendix B Analytic Solutions for Affine Models
$\quad-\kappa_{Y}\mathbf{C}(t)\bm{\theta}_{Y}.$ (27)
@@
$\frac{dg(t)}{dt}$ $=(1-\gamma)r-\frac{1-\gamma}{2\gamma}\operatorname{Tr}\left((\mathbf{A}^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})\right)-\frac{1-\gamma}{\gamma}\mathbf{B}(t)^{\top}(\bm{\sigma}_{Y}^{\top}\rho^{\top}\bm{\sigma})\Sigma^{-1}(\bm{\sigma}\mathbf{A})$
@@
$\quad+\mathbf{B}(t)^{\top}\kappa_{Y}\bm{\theta}_{Y}\quad+\frac{1}{2}\operatorname{Tr}(\mathbf{C}(t)\bm{\sigma}_{Y}\Phi\bm{\sigma}_{Y}^{\top}).$
@@
Appendix C Complete and Rigorous Proofs for the Policy‑Gap Bound (Theorem 1)
This appendix provides the complete, self‑contained analytic backbone for Theorem 1. All proof sketches have been expanded into detailed arguments, with explicit reliance on parabolic maximal‑regularity and monotone‑operator theory.
C.1 Table of Symbols
Symbol Meaning
$E_{v},E_{h},E_{f}$ Gradient / coefficient / source errors
$\varepsilon$ FOC residual norm
$\delta_{\mathrm{BPTT}}$ Costate estimation error (Def. 1)
$\underline{\Lambda}$ Lower eigenvalue from Lemma 4
$C_{S},C_{v}^{\prime\prime}$ PDE stability / coupled‑error constants
$L_{F},C_{\text{tot}}$ Final error amplification constants
C.2 Preliminaries and Notation
Throughout, we set the primary costate
$\lambda:=\nabla_{x}V^{\pi,c}.$
For any admissible control $(\pi,c)$ the sub‑value function $V^{\pi,c}$ solves
$-\partial_{t}V^{\pi,c}=\mathcal{L}_{\pi,c}[V^{\pi,c}]+f_{\pi,c},\qquad V^{\pi,c}(T,\cdot)=g,$
with oblique‑derivative boundary $\partial_{\mathbf{n}}V=0$ on $\partial\Omega$ to match the Skorokhod reflection (Lieberman, , 1996). Error quantities on $Q_{T}:=(0,T)\times\Omega$:
$h$ $:=(a_{\pi,c}-a_{*},\,b_{\pi,c}-b_{*}),$ $\quad E_{h}$ $:=\|h\|_{L^{q,p}},$
$E_{f}$ $:=\bigl{\|}e^{-\delta t}(U(c^{*})-U(c))\bigr{\|}_{L^{q,p}},$ $\quad E_{v}$ $:=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}},$
$\varepsilon$ $:=\|\mathcal{R}_{\mathrm{FOC}}\|_{L^{q,p}}.$
C.3 Minimal Assumptions
Assumption 6 (Baseline Well‑Posedness).
Fix $T>0$, a $C^{1,1}$ domain $\Omega$, mixed‑norm exponents $p>d+2,\;q\geq 2$ and a compact cylinder
$\mathcal{D}:=[0,T]\times[\underline{x},\bar{x}]\times\{\|y\|\leq R_{Y}\}.$
There exist positive constants $\underline{\nu},\overline{\nu},L_{\mathrm{coef}},C_{\mathrm{MR}},\eta_{0}$ such that
  1. 1.
    Reflected SDE. The Skorokhod‑reflected SDE with inward normal $\mathbf{n}\in C^{1,\alpha}$ is well posed.
  2. 2.
    Uniform ellipticity & VMO coefficients. $\underline{\nu}|\xi|^{2}\!\leq\!\xi^{\top}a_{\pi,c}\xi\!\leq\!\overline{\nu}|\xi|^{2}$ and $(a_{\pi,c},b_{\pi,c})$ have VMO/BMO bounds.
  3. 3.
    VMO‑smallness. On every cylinder $Q_{r}\subset\mathcal{D}$ with $r\leq R_{0}$ the VMO modulus is $\leq\eta_{0}<\bar{\eta}(d,p,q)$ of Dong and Kim, (2009).
  4. 4.
    Costate band & utility. $\underline{\lambda}\leq\lambda\leq\bar{\lambda}$ on $\mathcal{D}$ and $U(c)=c^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$.
C.4 Automatic Properties of the Model
Lemma 4 (Strong concavity of the Hamiltonian).
@@ C.4 Automatic Properties of the Model
($-\nabla^{2}_{(\pi,c)}H(t,x,y,\lambda,\eta,\pi,c)\;\succeq\;\underline{\Lambda}I_{n+1},$
where $\underline{\Lambda}>0$ depends only on $(a,b,\gamma,\Pi_{\max},\underline{\lambda},\bar{\lambda})$.
@@
HARA implies $U^{\prime\prime}(c)\leq-\kappa_{0}<0$ on the compact $c$‑interval, while the quadratic portfolio term contributes $-x^{2}\lambda\,\sigma^{\top}\Sigma\sigma\preceq-\kappa_{1}I_{n}$ since $x\in[\underline{x},\bar{x}]$ and $\Sigma\succ 0$. Summing gives the claim. ∎
Lemma 5 (Maximal regularity for $V^{\pi,c}$).
Under Assumption 6 and the VMO‑smallness $\eta_{0}<\bar{\eta}(d,p,q)$, both $V^{\pi,c}$ and $V^{*}$ lie in $W^{2,1}_{q,p}(Q_{T})$ with
$\|D^{2}V^{\pi,c}\|_{L^{q,p}}+\|\partial_{t}V^{\pi,c}\|_{L^{q,p}}\;\leq\;C_{\mathrm{MR}}\bigl{(}\|f_{\pi,c}\|_{L^{q,p}}+\|g\|_{W^{2-2/p}_{q}}\bigr{)}.$